# The Importance of Being Discrete: Measuring the Impact of Discretization in End-to-End Differentially Private Synthetic Data

### v2

@@ 5 Experimental Evaluation
In this section, we present our experimental evaluation aimed at addressing the four research questions introduced in Section 1.
@@ 5.1 RQ1: Discretization and Capturing Controlled Distributions
We begin by discussing the discretization of controlled distributions and use US1 (cf. Section 3.5) to measure how much utility is preserved.
Motivating Examples. Before diving into the details of our experiments, we first illustrate, visually, the effects of discretization on the six controlled distributions – see Figure 4. The differences between discretization strategies are clear in how they partition the space, particularly in the placement and spacing of bin edges, as well as the distribution of data across bins.
In Figure 16 and 17 in Appendix A, we also present two examples that visually highlight the importance of discretization in simple clustering and classification tasks. Overall, we observe that discretization can enhance modeling in certain scenarios.
Discretizers Utility (US1). Next, we measure how well the four discretizers capture the controlled distributions as per US1 (cf. Figure 3(a)). In Figure 5, we report the aggregated utility averaged across the datasets. Utility improves as the number of bins increases for all the discretizers. Remarkably, they approach the maximum achievable score (within 5%) when 250 bins are used, even under tight privacy constraints ($\epsilon=0.1$). This suggests that increasing the number of bins effectively offsets the added noise introduced by DP, likely because finer granularity reduces the distance between sampled points and their original values. At $\epsilon=\infty$, k-means is the best-performing discretizer, although it experiences the biggest decline when DP is applied. By contrast, while uniform remains unchanged with DP, PrivTree also shows minimal change in performance.
@@ 5.2 RQ2: Discretization and DP Generative Models
Next, we examine the four discretizers in combination with baseline DP generative modeling controlled distributions (US2; see Figure 3(b)) to understand their behavior in controlled settings. We then assess their impact when used with DP generative models on real datasets (US3, see Figure 3(c)).
Discretizers Utility (US2 vs. US1). As per US2, we adopt a simple DP generative model pipeline and compare its performance with US1, which only discretized the data and then sampled back. We plot an aggregated comparison between US2 and US1 for $\epsilon=0.1$ in Figure 6, and a more detailed view of US2 in Figure 7, showing the four discretizers’ behavior across a range of bins and two privacy budgets. Unlike US1, in US2, increasing the number of bins only helps to a certain point; using more bins beyond it actually harms utility (see Figure 6). This occurs because, for a fixed privacy budget, the noise allocation must be spread across all bins when modeling the distribution. As the number of bins increases over a certain point, the noise added per bin grows, ultimately degrading the captured distribution and reducing utility.
By comparing the discretizers (see Figure 7), we note that, with $\epsilon=\infty$, which closely mirrors US1, all discretizers continue to improve as the number of bins increases. However, under DP constraints, most discretizers either plateau or show a slight decline in utility, except for PrivTree, which incorporates an internal mechanism to limit the number of bins as $\epsilon$ decreases. Overall, PrivTree demonstrates the most promising performance (except for k-means, which outperforms it with up to 25 bins for $\epsilon=\infty$).
Figure 8: Utility of DP generative models and DP discretizers, Gas dataset (US3).
@@
Finally, for models with relatively low variability (i.e., PrivBayes, MST, and AIM), we observe a clear trend related to the number of bins. At lower epsilon values, the best results are achieved with fewer bins (darker dots are positioned lower). As epsilon increases, higher numbers of bins yield better results (darker dots shift upward). We further explore the importance of finding the optimal number of bins in RQ3 in Section 5.3.
@@
Optimal Discretizer (US3). As already reported in Figure 2 (in Section 1), optimizing the choice of discretizer (and the number of bins) can increase the utility of all DP generative models by 9.28% to 43.54% compared to the baseline choice, uniform with 20 bins. In all cases where the best results are achieved, PrivTree emerges as the optimal choice. Moreover, PrivTree consistently outperforms the other discretizers across various factors, namely, privacy budgets, number of bins, generative models, and datasets, achieving the highest performance in 63% of the cases, while k-means ranks second with 16%. This confirms previous claims that PrivTree is better than uniform [80].
PrivTree’s superiority is also visible from Figure 1, where the four discretizers are compared for $\epsilon=1$ (selected to ease presentation, though the same pattern holds for other $\epsilon$ values). Notably, the overall trend, i.e., increasing the number of bins improves utility up to a point, is consistent with the findings in US2 (cf. Figure 7). As for variability, PrivTree also achieves the best overall score, albeit by a small margin – more precisely, 0.05 compared to 0.06 for the other discretizers (cf. Figure 8).
@@
Figure 9: Utility of sampling strategies, averaged across four DP discretizers, six DP generative models, and three datasets (US3).
@@
Discretizers Sampling (US3). In Figure 9, we compare the two sampling methods – uniform and mixture. Both sampling methods can be applied to all discretizers in the post-processing step, when converting bin indices back to the original numerical domain. When using mixture sampling, we allocate 50% of the discretizer’s budget to fitting the truncated normal distributions of the bins.
@@
Using mixture sampling tends to outperform uniform under certain conditions – high epsilons and low number of bins. This is perhaps expected, as modeling the distribution of the values within a given bin adds to the expressiveness of the bin representation but at the cost of privacy. However, once we have enough bins (more than 10), reducing the discretization budget results in uniform sampling outperforming mixture. This is likely due to inaccurate modeling of mean and standard deviation per bin since the available budget needs to be distributed to too many bins. Since we tend to achieve the best results with the number of bins more than 10 (see Figure 1 and Figure 7), we use the uniform sampling by default.
Computational Performance (US3). We also compare the computational overhead of the discretizers and generative models with varying number of bins, reporting runtimes in Figure 10 and Figure 11, respectively. In terms of discretizers’ performance, quantile and k-means take longer to fit as the number of bins increases, e.g., from 2-10 secs for 5 bins to 40-100 secs for 250 bins. By contrast, both uniform and PrivTree are much faster (under one sec) and, more importantly, remain relatively unaffected by bin count, making them highly scalable.
@@
Take-Aways. Our experiments show that DP generative models are highly sensitive to the choice of discretization strategy and bin count, and optimizing them can increase utility by around 30% on average. While increasing the number of bins initially improves utility, doing so beyond a certain point starts to negatively impact it. We also find that PrivTree consistently outperforms other discretizers in terms of utility, variability, and computational efficiency.
@@ 5.3 RQ3: Discretization and Optimal Number of Bins
Optimized Strategy for Number of Bins (US2). Figure 12, we report the average utility of four strategies for automatically selecting the number of bins – $b^{\mathbf{DF}}$, $b^{\mathbf{RL}}$, $b^{\mathbf{FDR}}$, and $b^{\mathbf{SHI}}$ (introduced in Section 3.1) – alongside the number of bins that led to achieving the best utility for the same setting on controlled distributions (US2). The privacy budget is a key factor in determining the optimal number of bins; as the available budget increases, using more bins leads to higher utility – a trend already noted in Section 5.2. Among the four bin selection strategies, $b^{\mathbf{SHI}}$ performs best on average. However, implementing this strategy in a DP manner (i.e., ‘DPfying’ the formula) poses practical challenges, as it relies on numerous unperturbed queries to the data. Making all these queries DP would likely result in either inaccurate modeling or a ‘wasted’ privacy budget. Similarly, $b^{\mathbf{DF}}$ and $b^{\mathbf{FDR}}$ encounter the same limitations. By contrast, $b^{\mathbf{RL}}$ provides a better balance between simplicity and performance.
@@
Therefore, we select $b^{\mathbf{RL}}$ and attempt to optimize it for DP settings because it is data-independent, making it DP without any additional detriment to privacy. The optimized version of $b^{\mathbf{RL}}$, denoted as $b^{\mathbf{RL-OPT}}$, takes into account the privacy budget to reduce the number of bins and is introduced in Section 4.
@@
Optimal Number of Bins (US3). In Figure 13, we report the average utility for all tested bin counts and selection strategies. While $b^{\mathbf{RL-OPT}}$ yields the highest utility, achieving a 41% normalized improvement between the default (uniform with 20 bins) and the best scores, its absolute improvement over other methods remains modest. This leaves room for future work to develop improved strategies, possibly tailored to specific discretizers.
@@
Nevertheless, using $b^{\mathbf{RL-OPT}}$ bins for discretization and training a generative model saves privacy budget and reduces computational resources by eliminating the need for extensive hyperparameter search, which is required to achieve the best possible score. By contrast, hyperparameter searches generally require additional privacy budget, making them costly in DP settings [65, 37].
@@
Take-Aways. Our bin selection strategy, $b^{\mathbf{RL-OPT}}$, outperforms all tested bin counts and alternative strategies (and we believe there is actually potential for further improvement).
@@
Figure 15: Privacy leakage with provided and extracted domain (w/ and w/o DP) for the four DP discretizers (D) and two DP generative models (G) on a target record outside/inside the domain of the remaining data, Wine dataset (PS1)
5.4 RQ4: Discretization and DP Domain Extraction
Finally, we evaluate the impact of DP domain extraction on utility (US3) and privacy (PS1; see Figure 3(d)) across various discretizers and generative models applied to real datasets. We compare three scenarios: 1) the domain is assumed to be provided and set to that of the full dataset, 2) the domain is extracted directly from the input dataset (without DP), as done in numerous publicly available implementations and libraries [96, 67, 87, 51, 52, 45, 69, 12], and 3) the domain is extracted from the input dataset with DP. In the first two cases, the entire discretization privacy budget is allocated to the discretizer, while in the third, the budget is split evenly, with 50% used for domain extraction and 50% for discretization.
Utility (US3). In Figure 14, we present the average utility across the three domain extraction scenarios for each of the four discretizers, averaged across all models and datasets. The first two scenarios yield identical results (bars with lines/crosses), as the input dataset matches the full dataset. Introducing DP domain extraction (bars with circles) reduces utility across all discretizers – approximately 2.5% for uniform, quantile, and k-means, and 8% for PrivTree. Despite this drop, PrivTree still outperforms the others, even when they extract the domain directly without DP. This underscores the sensitivity of all discretizers to domain specification and highlights the need for future research in this area.
@@ 5.4 RQ4: Discretization and DP Domain Extraction
Privacy (PS1). Figure 15(a) illustrates the impact of the domain on a popular membership inference attack (MIA), GroundHog [76], across different discretizers and generative models. We train the generative models with $\epsilon=1$, and use $\epsilon=1$ for discretization (split in half between domain extraction and discretization for the third setting). Notably, the selected target record represents a worst-case scenario on a real dataset, which is permitted under DP assumptions and consistent with prior work [76, 4, 22]. The target contains two columns with values significantly larger than those of the remaining records (289 and 440 vs. 146.5 and 366.5).
@@
Extracting the domain directly from the input data (bars with horizontal lines), particularly the minimum and maximum of numerical columns, without a proper privacy mechanism not only breaks end-to-end DP guarantees but also exposes highly informative features to the adversary. As a result, the adversary achieves near-perfect performance in all cases, deeming the synthetic data non-private regardless of the discretizer or generative model used. This is equivalent to the default strategy (grey bars), which extracts the domain directly and applies uniform discretization.
@@
By contrast, adopting methods that respect the end-to-end DP pipeline, i.e., either by assuming the domain is provided (bars with crosses) or extracting it with DP (bars with circles), substantially reduces the adversary’s success rate. In all cases, with a single exception (k-means with a provided domain), the attack success rate drops to the level of random guessing. This has several important implications.
@@
First, extracting the domain in a DP-compliant way is sufficient to protect against the specific adversary used in the GroundHog attack [76]. Furthermore, the adversary’s success remains low even in settings that could be considered non-private, e.g., (discretizer, generator)-$\epsilon$ values of (100, 100), as shown in Figure 15(b) (again, which the exception of k-means). Similar results are observed for (1, 100), (1, 100), and even (1, 1,000), though not shown due to space constraints. The attack only becomes more effective at higher discretizer values, like (1,000, 1) and (1,000, 1,000) (see Figure 15(c)); also, recall that it achieves nearly 100% success when the domain is directly extracted from the data. This suggests that the effectiveness of the GroundHog attack may stem primarily from domain leakage, rather than intrinsic vulnerabilities in the generative model itself. [^3] These findings are specific to GroundHog [76]; domain extraction’s impact on attacks tailored to PrivBayes, MST, or other MIAs remains unclear and left to future work. To further validate this, we conduct an additional experiment on a target record located farther away from the others but still within their domain (see Figure 15(d)) and observe that the attack’s success remains near random across all settings.
@@
Second, adopting a DP domain extraction strategy can help address privacy vulnerabilities identified by prior research [4, 22] in popular model implementations and libraries [67, 69] that extract the domain directly from the input data. In other words, incorporating such techniques can make DP generative model implementations more robust and better align them with end-to-end DP guarantees.
@@
Finally, while extracting the domain in a DP way slightly reduces the adversary’s success rate compared to using a provided domain (for $\epsilon\leq 100$), this may come at the cost of utility as already shown in Figure 14. Practitioners should therefore prefer using a trusted, externally provided domain (e.g., codebooks for census data) when available, rather than spending additional privacy budget to infer it. However, further research is needed to explore these trade-offs and to develop more effective methods for DP domain extraction.

### v4

@@ 5 Experimental Evaluation
In this section, we present our experimental evaluation aimed at addressing the four research questions introduced in Section˜1.
@@ 5.1 RQ1: Discretization and Capturing Controlled Distributions
We begin by discussing the discretization of controlled distributions in US1 (cf. Section˜3.5) and measure how much utility is preserved.
Motivating Examples. Before diving into the details of our experiments, we first illustrate, visually, the effects of discretization on the six controlled distributions – see Figure˜4. The differences between discretization strategies are clear in how they partition the space, particularly in the placement and spacing of bin edges, as well as the distribution of data across bins.
In Figure˜16 and 17 in Appendix A, we also present two examples that visually highlight the importance of discretization in simple clustering and classification tasks. Overall, we observe that discretization can enhance modeling in certain scenarios.
Discretizers Utility (US1). Next, we measure how well the four discretizers capture the controlled distributions as per US1 (cf. Figure˜3(a)). In Figure˜5, we report the aggregated utility averaged across the datasets. Utility improves as the number of bins increases for all discretizers. Remarkably, they approach the maximum achievable score (within 5%) when 250 bins are used, even under tight privacy constraints ($\epsilon=0.1$). This suggests that increasing the number of bins effectively offsets the added noise introduced by DP, likely because finer granularity reduces the distance between sampled points and their original values. At $\epsilon=\infty$, k-means is the best-performing discretizer, although it experiences the biggest decline when DP is applied. By contrast, while uniform remains unchanged with DP, PrivTree also shows minimal change in performance.
@@ 5.2 RQ2: Discretization and DP Generative Models
Next, we examine the four discretizers in combination with baseline DP generative modeling controlled one-dimensional distributions (US2; see Figure˜3(b)) to understand their behavior in controlled settings. We then assess their impact when used with DP generative models on real high-dimensional datasets (US3, see Figure˜3(c)).
@@
Discretizers Utility (US2 vs. US1). As per US2, we adopt a simple DP generative model pipeline and compare its performance with US1, which only discretized the data and then sampled back. We plot an aggregated comparison between US2 and US1 for $\epsilon=0.1$ in Figure˜6, and a more detailed view of US2 in Figure˜7, showing the four discretizers’ behavior across a range of bins and two privacy budgets. Unlike US1, in US2, increasing the number of bins only helps to a certain point; using more bins beyond it actually harms utility (see Figure˜6). This occurs because, for a fixed privacy budget, the noise allocation must be spread across all bins when modeling the distribution. As the number of bins increases over a certain point, the noise added per bin grows, ultimately degrading the captured distribution and reducing utility. This trend generalizes to other $\epsilon$ values as well, with higher budgets improving overall utility and shifting the inversion point to a larger number of bins (for ease of presentation, we only visualize the trend for $\epsilon=0.1$).
By comparing the discretizers (see Figure˜7), we note that, with $\epsilon=\infty$, which closely mirrors US1, all discretizers continue to improve as the number of bins increases. However, under DP constraints, most discretizers either plateau or show a slight decline in utility, except for PrivTree, which incorporates an internal mechanism to limit the number of bins as $\epsilon$ decreases. Overall, PrivTree demonstrates the most promising performance (except for k-means, which outperforms it with up to 25 bins for $\epsilon=\infty$).
Finally, in Figure˜18 in Appendix B, we present the same information as in Figure˜7 but disaggregate the three utility metrics in US2. Although the metric ranges vary, the discretizers show consistent behavior across all metrics, ensuring that each metric is well represented in the final aggregation.
Generative Models Utility (US3). Figure˜8 reports the utility results on the Gas dataset, broken down by the four discretizers and six generative models, as per US3 (numbers in boxes/parentheses indicate standard deviation). All generative models are trained with a privacy budget $\epsilon$ in the range {0.01, 0.1, 1, 10, 100, $\infty$}, except for AIM, which is limited to $\epsilon\leq 1$ due to computational constraints (it takes over 60 mins for $\epsilon\geq 10$). We observe that both the choice of discretizer and the number of bins have a significant impact on the performance of all generative models.
@@
Finally, for models with relatively low variability (i.e., PrivBayes, MST, and AIM), we observe a clear trend related to the number of bins. At lower epsilon values, the best results are achieved with fewer bins (darker dots are positioned lower). As epsilon increases, higher numbers of bins yield better results (darker dots shift upward). We further explore the importance of finding the optimal number of bins in RQ3 in Section˜5.3.
@@
Optimal Discretizer (US3). As already reported in Figure˜2 (in Section˜1), optimizing the choice of discretizer (and the number of bins) can increase the utility of all DP generative models by 9.28% to 43.54% compared to the baseline choice, uniform with 20 bins. In all cases where the best results are achieved, PrivTree emerges as the optimal choice. Moreover, PrivTree consistently outperforms the other discretizers across various factors, namely, privacy budgets, number of bins, generative models, and datasets, achieving the highest performance in 63% of the cases, while k-means ranks second with 16%. This confirms previous claims that PrivTree is better than uniform [81].
@@
Figure 9: Utility of sampling strategies, averaged across four DP discretizers, six DP models, three datasets (US3).
PrivTree’s superiority is also visible from Figure˜2, where the four discretizers are compared for $\epsilon=1$ (selected to ease presentation, though the same pattern holds for other $\epsilon$ values). Notably, the overall trend, i.e., increasing the number of bins improves utility up to a point, is consistent with the findings in US2 (cf. Figure˜7). As for variability, PrivTree also achieves the best overall score, albeit by a small margin – more precisely, 0.05 compared to 0.06 for the other discretizers (cf. Figure˜8).
@@
Discretizers Sampling (US3). In Figure˜9, we compare the two sampling strategies – uniform and mixture. Both sampling methods can be applied to all discretizers in the post-processing step, when converting bin indices back to the original numerical domain. When using mixture sampling, we allocate 50% of the discretizer’s budget to fitting the truncated normal distributions of the bins.
@@
Using mixture sampling tends to outperform uniform under certain conditions – high epsilons and low number of bins. This is perhaps expected, as modeling the distribution of the values within a given bin adds to the expressiveness of the bin representation but at the cost of privacy. However, once we have enough bins (more than 10), reducing the discretization budget results in uniform sampling outperforming mixture. This is likely due to inaccurate modeling of mean and standard deviation per bin since the available budget needs to be distributed to too many bins. Since we tend to achieve the best results with the number of bins more than 10 (see Figure˜2 and Figure˜7), we use the uniform sampling by default.
@@
Computational Performance (US3). We also compare the computational overhead of discretizers and generative models with varying number of bins, reporting runtimes in Figure˜10 and Figure˜11, respectively. In terms of discretizers’ performance, quantile and k-means take longer to fit as the bin count increases, e.g., from 2-10 secs for 5 bins to 40-100 secs for 250 bins. By contrast, uniform and PrivTree are much faster (under a sec) and, importantly, remain relatively unaffected by bin count, making them highly scalable.
Take-Aways. Our experiments show that DP generative models are highly sensitive to the choice of discretization strategy and bin count, and optimizing them can increase utility by ~30% on average. While increasing the number of bins initially improves utility, doing so beyond a certain point starts to negatively impact it. We also find that PrivTree consistently outperforms other discretizers in terms of utility, variability, and computational efficiency.
@@ 5.3 RQ3: Discretization and Optimal Number of Bins
Optimized Strategy for Number of Bins (US2). In Figure˜12, we report the average utility of four strategies for automatically selecting the number of bins – $b^{\mathbf{DF}}$, $b^{\mathbf{RL}}$, $b^{\mathbf{FDR}}$, and $b^{\mathbf{SHI}}$ (introduced in Section˜3.1) – alongside the number of bins that led to achieving the best utility for the same setting on controlled distributions (US2). The privacy budget is a key factor in determining the optimal number of bins; as the budget increases, using more bins leads to higher utility – a trend already noted in Section˜5.2. Among the four bin selection strategies, $b^{\mathbf{SHI}}$ performs best on average. However, implementing this strategy in a DP manner (i.e., ‘DPfying’ the formula) poses practical challenges, as it relies on numerous unperturbed queries to the data. Making these queries DP would likely result in either inaccurate modeling or ‘wasted’ privacy budget. Similarly, $b^{\mathbf{DF}}$ and $b^{\mathbf{FDR}}$ encounter the same limitations. By contrast, $b^{\mathbf{RL}}$ provides a better balance between simplicity and performance.
@@
Therefore, we select $b^{\mathbf{RL}}$ and attempt to optimize it for DP settings because it is data-independent, making it DP without any additional detriment to privacy. Our optimized version of $b^{\mathbf{RL}}$, denoted as $b^{\mathbf{RL-OPT}}$, takes into account the privacy budget to reduce the number of bins and is introduced in Section˜4 (we also explored more complex correction terms but none yielded better results).
@@
Optimal Number of Bins (US3). In Figure˜13, we report the average utility for all tested bin counts and selection strategies. While $b^{\mathbf{RL-OPT}}$ yields the highest utility among all strategies, achieving a 41% normalized improvement between the default and the best achievable score via exhaustive search, its absolute improvement remains modest. This leaves room for future work to develop improved strategies, possibly tailored to specific discretizers.
@@
Nevertheless, using $b^{\mathbf{RL-OPT}}$ bins for discretization and training a generative model saves privacy budget and reduces computational resources by eliminating the need for extensive hyperparameter search, which is required to achieve the best score. By contrast, hyperparameter searches generally require additional privacy budget, making them costly in DP settings [66, 38].
@@
Take-Aways. Our optimized bin selection strategy, $b^{\mathbf{RL-OPT}}$, outperforms all tested bin counts and alternative strategies (and we believe there is actually potential for further improvement).
5.4 RQ4: Discretization and DP Domain Extraction
Finally, we evaluate the impact of DP domain extraction on utility (US3) and privacy (PS1; see Figure˜3(d)) across various discretizers and generative models applied to real datasets. We compare three scenarios: 1) the domain is assumed to be provided and set to that of the full dataset, 2) the domain is extracted directly from the input dataset (without DP), as done in numerous publicly available implementations and libraries [97, 68, 88, 52, 53, 46, 70, 12], and 3) the domain is extracted from the input dataset with DP. In the first two cases, the entire discretization privacy budget is allocated to the discretizer, while in the third, the budget is split evenly, with 50% used for domain extraction and 50% for discretization.
Utility (US3). In Figure˜14, we present the average utility across the three domain extraction scenarios for each of the four discretizers, averaged across all models and datasets. The first two scenarios yield identical results (bars with lines/crosses), as the input dataset matches the full dataset. Introducing DP domain extraction (bars with circles) reduces utility across all discretizers – approximately 2.5% for uniform, quantile, and k-means, and 8% for PrivTree. Despite this drop, PrivTree still outperforms the others, even when they extract the domain directly without DP. This experiment underscores the sensitivity of all discretizers to domain specification and highlights the need for future research in this area.
@@ 5.4 RQ4: Discretization and DP Domain Extraction
Figure 15: Privacy leakage with provided and extracted domain (w/ and w/o DP) for four DP discretizers (D) and two DP models (G) on a target record outside/inside the domain of the remaining data, GroundHog [77], Wine dataset (PS1).
@@
Privacy (PS1). Figure˜15(a) illustrates the impact of the domain on a popular membership inference attack (MIA), GroundHog [77], across different discretizers and generative models. [^3] We repeat the experiment using Querybased [33] and observe very similar results; due to space limitation, visualizations are provided in Figure 19 in Appendix C. We train the generative models with $\epsilon=1$, and use $\epsilon=1$ for discretization (split in half between domain extraction and discretization for the third setting). Notably, the selected target record represents a worst-case scenario on a real dataset, which is permitted under DP assumptions and consistent with prior work [77, 4, 22]. The target contains two columns with values significantly larger than those of the remaining records (289 and 440 vs. 146.5 and 366.5).
@@
Extracting the domain directly from the input data (bars with horizontal lines), particularly the minimum and maximum of numerical columns, without a privacy mechanism not only breaks end-to-end DP guarantees but also exposes highly informative features to the adversary (either directly obtained by GroundHog [77] or indirectly via queries by Querybased [33]). As a result, the adversary achieves near-perfect performance in all cases, deeming the synthetic data non-private regardless of the discretizer or generative model used. This is equivalent to the default strategy (grey bars), which extracts the domain directly and applies uniform discretization.
@@
By contrast, adopting methods that respect the end-to-end DP pipeline, i.e., either by assuming the domain is provided (bars with crosses) or extracting it with DP (bars with circles), substantially reduces the adversary’s success rate. In all cases, with a single exception (k-means with a provided domain), the attack success rate drops to the level of random guessing. This has several implications.
@@
First, extracting the domain in a DP-compliant way is sufficient to protect against the specific adversaries used in GroundHog and Querybased attacks. Furthermore, the adversaries’ success remains low even in settings that could be considered non-private, e.g., (discretizer, generator)-$\epsilon$ values of (100, 100), as shown in Figure˜15(b) (again, which the exception of k-means). Similar results are observed for (1, 100), (1, 100), and even (1, 1,000), though not shown due to space constraints. The attack only becomes more effective at higher discretizer values, like (1,000, 1) and (1,000, 1,000) (see Figure˜15(c)); also, recall that it achieves nearly 100% success when the domain is directly extracted from the data. This suggests that the effectiveness of the GroundHog/Querybased attacks may stem primarily from domain leakage, rather than intrinsic vulnerabilities in the generative model itself. [^4] These findings are specific to GroundHog [77] and Querybased [33], which are among the state-of-the-art MIAs on synthetic data; domain extraction’s impact on attacks tailored to PrivBayes, MST, or other MIAs remains unclear and left to future work. To further validate this, we conduct an additional experiment on a target record located farther away from the others but still within their domain (see Figure˜15(d) and Figure˜19(d) in Appendix C) and observe that the attack’s success remains near random across all settings.
@@
Second, adopting a DP domain extraction strategy can help address privacy vulnerabilities identified by prior research [4, 22] in popular model implementations/libraries [68, 70] that extract the domain directly from the input data. In other words, incorporating such techniques can make DP generative model implementations more robust and better align them with end-to-end DP guarantees.
@@
Finally, while extracting the domain in a DP way slightly reduces the adversary’s success rate compared to using a provided domain (for $\epsilon\leq 100$), this may come at the cost of utility as already shown in Figure˜14. Practitioners should therefore prefer using a trusted, externally provided domain (e.g., codebooks for census data) when available, rather than spending additional privacy budget to infer it. However, further research is needed to explore these tradeoffs and to develop more effective methods for DP domain extraction.